OpenStudy (natsent06):

c=5/9(f-32) what if its 45℃ how cold is it in ℉

OpenStudy (chosenmatt):

@WolfMaster

OpenStudy (natsent06):

wut who medaled u

OpenStudy (natsent06):

@iGreen

OpenStudy (natsent06):

@iGreen

OpenStudy (natsent06):

can u help @iGreen

OpenStudy (natsent06):

omg @zBrandz23 ur pic is ADORABLE!!

OpenStudy (igreen):

Just plug in 45 for C in the equation and solve for F: c=5/9(f-32) 45=5/9(f-32) Distribute 5/9 into the parenthesis: 45 = 5/9f - 17.78 Add 17.78 to both sides: 62.78 = 5/9f Divide 5/9 to both sides or multiply by its reciprocal, 9/5. What's 9/5 * 62.78? @natsent06

OpenStudy (igreen):

Oh nevermind..you changed it.. xD

OpenStudy (natsent06):

what do u mean i changed it

OpenStudy (zbrandz23):

Thanks @natsent06 out of all the wolf puppy pics this pic was most adorable I could find:D

OpenStudy (zbrandz23):

He's talking about my pic @natsent06

OpenStudy (igreen):

He had a different picture before and I made it for him..lol.

OpenStudy (zbrandz23):

It was a zrbra owl with dark aviators

OpenStudy (natsent06):

113.004?

OpenStudy (zbrandz23):

zebra